### v1

# LLM-enhanced Scene Graph Learning for Household Rearrangement

Abstract.
The household rearrangement task involves spotting misplaced objects in a scene and accommodate them with proper places. It depends both on common-sense knowledge on the objective side and human user preference on the subjective side. In achieving such task, we propose to mine object functionality with user preference alignment directly from the scene itself, without relying on human intervention. To do so, we work with scene graph representation and propose LLM-enhanced scene graph learning which transforms the input scene graph into an affordance-enhanced graph (AEG) with information-enhanced nodes and newly discovered edges (relations). In AEG, the nodes corresponding to the receptacle objects are augmented with context-induced affordance which encodes what kind of carriable objects can be placed on it. New edges are discovered with newly discovered non-local relations. With AEG, we perform task planning for scene rearrangement by detecting misplaced carriables and determining a proper placement for each of them. We test our method by implementing a tiding robot in simulator and perform evaluation on a new benchmark we build. Extensive evaluations demonstrate that our method achieves state-of-the-art performance on misplacement detection and the following rearrangement planning.
journal: TOGcopyright: noneccs: Computer systems organization Robotics
1. Introduction
Scene analysis is a long-standing problem in computer graphics and vision (Patil et al., 2024) and finds many applications ranging from scene modeling/generation (Fisher et al., 2015) and VR/AR (Li et al., 2021) to various robotic applications (Zhang et al., 2023). Among the applications, some, e.g., robot-operated household rearrangement or room tiding, requires a much more in-depth analysis of the scenes than others (e.g., object goal navigation), due to the involvement of both common-sense knowledge on the objective side and human user preference on the subjective side.
When conducting a room tiding task, the agent needs to spot the misplaced objects and accommodate them in proper places (Sarch et al., 2022). This is quite challenging, especially in an unseen environment, as it relies on a fine-grained and deep understanding of the functionality of the objects. The determination of object functionality often requires a high-order analysis which relies on a larger context (Hu et al., 2018). Moreover, object functionality is often hinged on the house owner’s taste or preference, i.e., the owner’s intent on how an object is being used. For example, the owner could take a cabinet either as a closet or as a bookshelf, depending on where it is located and mainly what objects are placed in.
Large Language Models (LLMs) have been found extremely powerful in challenging, zero-shot task planning including scene rearrangement (Kant et al., 2022; Wu et al., 2023) due to the strong capability of common sense reasoning. However, since the built-in knowledge about object functionality and placement of an LLM is not grounded to the specific scene being considered, scene grounding is usually indispensable to ensure the executability of the generated plans and their relevance to the scene (Rana et al., 2023).
Another important issue to consider is that LLM ignores human preference totally if the prompt is not carefully designed or tuned. In the work of TidyBot (Wu et al., 2023), this is alleviated by explicitly injecting user preference as exemplar placements via prompts and having the LLM to summarize user preference out of the the exemplars. Note that the exemplar placements need to be specific to the scene in consideration to make the summarization scene-grounded, making the exemplar collection tedious. Han et al. (2024) propose an iterative self-training method to align the LLM planner with user preferences, which again relies on user feedback.
In this work, we propose to mine object functionality with user preference alignment directly from the scene itself, without relying on human intervention. With our method, the user instruction can be as simple as “tide the house” and the agent can conduct scene rearrangement automatically based on an in-depth analysis of object functionality and proper reasoning of user preference. One example is shown in Figure 1. Based on our deep analysis of personalized functionality, it appears the user would preferring utilizing the trunk as an unconventional surface for a table. Although this may deviate from the traditional use of a standard trunk, it aligns perfectly with the user’s unique desires for a more tailored and comfortable space.
To achieve that, we propose LLM-enhanced in-depth scene analysis to extract affordance for each carriable object in the scene reflecting simultaneously common-sense knowledge and personalized preference, both with scene grounding. To do so, we work with scene graphs which is an LLM-friendly representation for scene analysis. Starting with a vanilla scene graph, we harness LLM to enhance it into an affordance-enhanced graph (AEG) with information-enhanced nodes and newly discovered edges (relations). In particular, we construct a prompt containing the initial scene graph (in XML format) and the RGB-D observations for a multi-modal foundation model which we use GPT4V in this work. GPT4V then gradually converts the scene graph into an AEG where the nodes of receptacle objects are augmented with context-induced affordance mined by GPT4V. In particular, context-induced affordance encodes what kind of carriables can be placed on it. Some new edges are discovered based on analyzed non-local relations.
With AEG, we perform scene rearrangement in two steps. The first step is misplacement detection. For each carriable, we feed its context-induced affordance, its associated receptacle objects and the task instruction into GPT4V. GPT4V then outputs for each carriable a probability of misplacement. In the second step, we choose the top few most probably misplaced carriables and query GPT4V for potential receptacles for correct placement. A straightforward method for receptable query is to put all candidate receptacles in the scene into a very long prompt which may cause degraded accuracy and even hallucinations in output. We, instead, opt to pre-compile task-related affordance (affordance with task description) of all receptacles into an external memory in a similar spirit to Retrieval Augmented Generation (RAG). This way, the prompt for receptacle query is composed of a few most relevant entries retrieved from the memory, thus significantly improving query quality.
Figure 2. Given a scene graph (SG) with key frames, we utilize Large Language Model (LLM) to perform context-induced affordance analysis for all objects in the scene. These affordances are incorporated into the SG, updating the nodes and edges to construct an Affordance Enhanced Graph (AEG). We then evaluate the appropriateness of the current placements of all carriable objects in the AEG based on the affordance information of the objects and their receptacles, identifying misplaced items. For each misplaced carriable, we rate the suitability of each receptacle in the AEG as a placement target. The top k suitable receptacles are selected as candidates, and their affordances are retrieved as prompt to the LLM to generate the placement decision.
We test our method by implementing a robot system for indoor scene rearrangement in the Habitat 3.0 simulator (Puig et al., 2023). Entering an indoor scene, the robot first performs an RGB-D reconstruction and extracts a vanilla scene graph for the scene. It then transforms the scene graph into an AEG and plans actions for the rearrangement task automatically through prompting GPT4V with the scene graph and selected RGB keyframes. We also contribute a benchmark for evaluating the performance of scene rearrangement through annotating the Habitat Synthetic Scenes Dataset (HSSD 200) (Khanna et al., 2023) with ground-truth receptacles for each carriable. Extensive evaluations on existing benchmark and our new benchmark both demonstrate that our method achieves state-of-the-art performance on misplacement detection and the following rearrangement planning in the household rearrangement task.
Our main contributions include:
  • We introduce LLM-enhanced scene graph learning for deep mining of object functionality with user preference alignment, which results in an affordance-enhanced graph (AEG) encoding context-induced affordances of receptacles.
  • Based on AEG, we propose LLM-assisted misplacement detection and correct placement determination, with carefully designed mechanisms for improved efficiency.
  • We provide a benchmark of scene rearrangement by annotating a scene dataset with ground-truth receptacles. With the benchmark, we test our method with an end-to-end robot system for scene rearrangement implemented in simulator.
2. Related Works
Scene understanding and affordance analysis
Scene understanding is an important area of research that can be broadly categorized into perception tasks (Georgakis et al., 2017; Silberman et al., 2012; Zhang et al., 2020) and application tasks (Achlioptas et al., 2020; Azuma et al., 2022; Duan et al., 2022). As the field of embodied AI develops, scene understanding research has become more focused on understanding how scenes support the behavior of embodied agents such as human avatars and robots, i.e affordance (Gibson, 1977).
Affordance prediction has been approached at both the object level and the scene level. Object-level methods focuses on understanding how humans use items (Nagarajan et al., 2019; Koppula and Saxena, 2014; Fang et al., 2018; Deng et al., 2021; Nagarajan et al., 2020), while scene-level research emphasizes how to naturally place a human avatar within a scene (Wang et al., 2024; Li et al., 2019; Gupta et al., 2011; Kulal et al., 2023; Grabner et al., 2011; Zhu et al., 2016; Savva et al., 2014, 2016). Different from the aforementioned works which analyze affordance from a human interaction aspect, some studies focus on the analysis object-object affordance (Mo et al., 2022; Tang et al., 2021; LI et al., 2022). Their methods predict the physical interaction between objects, such as placement and stacking. (Tang et al., 2021; LI et al., 2022) extract potential object interactions within a given scene. However, all these object-object affordance works do not consider ¡scene affordance¿ (need to be defined in the introduction) of the objects, which requires an in-depth analysis of factors including the object’s geometry, pose, and the scene context, such as its position in the scene and the objects already interacting with it.
Foundation models for scene understanding
Visual language models (Radford et al., 2021) and large language models (Achiam et al., 2023) have been recently used in scene understanding tasks (Jatavallabhula et al., 2023; Hong et al., 2023; Wang et al., 2023). Some works(Li et al., 2023; Nguyen et al., 2023; Van Vo et al., 2023; Zhang et al., 2024) use or distill the rich commonsense provided by foundation models to identify affordance for a sole object. (Rana et al., 2023; Rajvanshi et al., 2023; Ni et al., 2023) use the scene graph to ground LLMs for navigation task. However, they just use the scene graph as a map without extracting object interaction information hiding in the scene graph. While in our work, we focus on using scene graph to ground LLMs in analyzing the object affordances, the analysis results are then sent back to the scene graph for deeper analysis.
Rearrangement
Rearrangement is a special embodied AI task to test the robots’ intelligence (Batra et al., 2020). In a given initial state of the scene, robots are tasked with manipulating the poses of objects to align with desired goal states. These goal states can be represented through various means such as object pose transformations(Liu et al., 2021), language instructions(Liu et al., 2022, 2023), or visual images (Weihs et al., 2021; Huang et al., 2023). It is worth noting that in certain cases like household tidying, robots may need to rearrange objects based on the original scene and their experiences. This necessitates the robot to have human-like common sense and the ability to decipher user preferences embedded within the original scenarios. Previous works in household rearrangement rearrange objects without considering the fine-grain affordances constrained by the scene (Kant et al., 2022; Sarch et al., 2022). In our method, we analyze the object affordance with scene graph and LLMs, which enable us rearrange the households while follow the original functionality of the object as much as possible.
3. Method
Figure 2 presents a overview of our approach. Starting with a 3D scene, abstracted into a scene graph where each object node is linked to a key frame, our method initiates by harnessing the contextual details of each node. This information is used to augment the graph, infusing it with induced affordances and uncovering new semantic connections between objects. Adopting the objective of ”tidying the house,” the affordance-enriched graph serves as the foundation for identifying objects that are out of place. To determine the most suitable location for each misplaced object, our system initially extracts a set of potential receptacle candidates from the graph. These candidates are then meticulously evaluated and compared to arrive at the optimal placement decision. It is important to highlight that all the pivotal steps of our methodology are executed in a zero-shot learning context, leveraging the advanced capabilities of the Large Language Model (LLM), specifically GPT4V.
Figure 3. Context-induced affordance analysis for objects in the scene graph. (a) For each object, we conduct a local analysis based on its textual contents related to its neighborhood in the graph and its key frame image, storing the results in the nodes as local affordances. (b) We construct object-area-room hierarchies and aggregate the local affordances of all objects within an area. Using LLM, we summarize the content and functionality of the entire area based on the aggregated information, obtaining descriptions for all areas within the room. (c) We use the descriptions of all areas in the room as the global context and assign this context to all nodes within the room to allow the LLM to identify meaningful semantic context information for each object. With this meaningful context, we construct new semantic edges between objects in the same room and update the previously analyzed local affordances, completing the context-induced affordance analysis of the scene graph.
Figure 4. Local affordance analysis of an example receptacle object.
3.1. Context-induced Affordance Analysis
In this section, we will explain how the initial scene graph is enhanced by using LLM to analyze the contextual information to in a local-to-global manner, as shown in Figure 3.
Initial scene graph
To simplify the problem setting, we assume that a vanilla scene graph $S$ has already been constructed from a RGBD scanning sequence for the given scene, where each node represents an object instance in the scene and two nodes will be connected with an edge if they are close enough. Moreover, several attributes of the object instance are stored in each node, including the instance ID, category label, located room label, and rearrangement type (carriable, receptacle, or others), where the rearrangement type follows the definitions from the Vocabulary Mobile Manipulation (OVMM)(Yenamandra et al., 2023) problem. In the Household task, only carriable objects can be rearranged, receptacle objects serve as targets for placing carriable objects, and other objects do not interact with the agent during the task but can provide contextual information for other objects. The relationship between two nearby objects is also categories into three different types, including near, on and support, and stored on the connecting edge. Note that if the scene graph is not given, it can be easily constructed following previous works, and details about the graph construction process can be found in the supplementary material.
To help with the inference, we further associate each node with a representative RGB frame selected from the input RGBD scanning sequence to provide a visual context. In more details, for each object instance, we first calculate the number of its projected pixels across all RGB-D images, creating an object-to-image-pixel matrix ($N*M$), where $N$ is the number of object instance and $M$ is the number of frames. To select the most informative frame, we identify all the neighboring objects and sum up their projected pixel numbers together with that of the current instance on each frame. The frame with the maximal number of projected pixels is selected, as it contained most visual information of surrounding context of the instance to be considered.
Local context analysis
Local context refers to the neighboring information stored in the scene graph $S$.
For each receptacle and other object in the $S$, we input all the textual contents related to its neighborhood and organize them into a descriptive text with the following template: “The name of this receptacle is [], its category is [], it is located in the room [], its relationships with surrounding objects are: it is near these objects: [], it supports these objects: [], it is supported by these objects: []. ” Note that if the given object does not have a certain type of edge, then the corresponding description will not be generated. Then we combine this template description with its associated key frame, and design a prompt for LLM to analyze the functionality of the object reflected in this scene via Chain of Thought (coT) reasoning. The final output consists of information related to Geometry & Position, Relationships, Functionality, and Fine-grained Category.
Figure 4 shows the local affordance analysis result of an example receptacle. In this example, the LLM first analyzes the geometric features of a table based on the visual information provided by the corresponding image. It combines this information with contextual information given by the scene graph to infer that the table is positioned against the wall at the entrance of the living room. Then, leveraging the position information and the supporting objects of the table in the scene graph, the LLM infers that the table can be likely used for placing outdoor wear. In the functionality module, based on the preceding analysis, the LLM further deduces that the table is suitable for placing items such as shoes that are typically found near entrances, as well as keys and other items that are temporarily stored when entering or leaving. As a conclusion, based on the reasoning above, the LLM assigns a new fine-grained category, ”Entryway organizing console”, to the table based on the above analysis.
Note that when conducting local context analysis, the information enhancement for carriable objects is a bit different, as its local context within the environment frequently changes, and thus we simplify the analysis without consider the position and relationship. For carriable objects, we input the category of object as well as its key frame image, requiring LLM to output the affordance of object with the following information: Geometry & Functionality and Fine-grained Category.
Global context analysis
Although our vanilla scene graph $S$ can effectively describe the surrounding context of a receptacle, the functionality of the receptacle may also depends on objects that are far away thus without connecting by an edge in $S$. For example, a sofa and a TV have a strong functional relationship (people sit on the sofa to watch TV), but they are typically arranged at a considerable distance from each other. Therefore, to consider more global context in the scene, we first build a hierarchical structure of $S$ and use it to bridge objects that are spatially distant yet functionally interrelated. In this way, new semantic edges are added, and the information stored in each object node is also updated, resulting in a new affordance-enhanced graph (AEG), as shown in Figure 3.
The details of the hierarchy construction process can be found in the supplementary material. Note that objects are first grouped into an intermediate granularity level called “area”, and then further merged into a room. Using this ”object-area-room” hierarchy to aggregate contextual information is more efficient than aggregating directly from objects to rooms. This method allows for layered, structured storage of aggregated information, preventing the loss of important details.
With the “object-area-room” hierarchy, we first use LLM to aggregate contextual information for each area separately. For each area, we collect the node information stored in $S_{l}$ of all the objects inside the area, and design a prompt for LLM to generate the following content based on a fixed template: Description and Fine-grained name. Therefore, each area will be associated with the above information together with the list of the object inside this area. We then integrate all the information from the areas within a room and use it as the global context for the room to extract distant context for all the receptacle and fixed objects within the room.
For each receptacle, we input the basic spatial information template description, its local affordance analysis as well as the global context of the room to the LLM. With this inputs and a designed prompt, LLM can extract the distant but meaningful context from the global information and add these contexts as semantic edges to the SG. These context output will be in a format of a list of objects, indicating objects within the room that is functionally related to the given receptacle, as well as the specific functional relationship between these objects and the given receptacle. After obtaining this meaningful context, we add semantic directed edges from the given receptacle to related objects, reflecting their functional relations.
In the end, we use LLM to update context-induced affordance based on the previous affordance analysis, the additional global context, as well as the semantic edges, with the same output format as in the local context afforadnce. In this way, our context-induced affordance includes both surrounding spatial context and distant semantic context, resulting in a more comprehensive understanding of receptacle’s functionality in this room. Note that we do not use global context to update the affordance of carriables for the same reason in local analysis. We refer the final scene graph with updated node information and newly added semantic edges as Affordance Enhanced Graph (AEG) and use it for further misplacemnet detection as well as rearrangement planning.
Figure 5. Misplace object scoring with the LLM.
Figure 6. We use a score-based retrieval-augmented method for placement decision generation. First, we construct a context-induced affordances database for all receptacles in the AEG. For each carriable object to be rearranged, we treat the object as a query and evaluate the suitability of each receptacle in the database as a placement target using the LLM scorer. The top k receptacles with the highest scores are selected as candidate receptacles. We retrieve their context-induced affordances from the database and use them along with the query as prompt input to the LLM to generate the best placement plan.
3.2. Misplaced Object Detection
Since our task is to “tide the house”, we are looking for carriables that currently have an unreasonable placement and need to be rearranged to complete the task. To achieve this, we introduced a LLM-based placement scorer to rate the suitability of a placement based on the affordance and description of carriable-receptacle pair. With this placement scorer, we can determine if a carriable object is appropriately placed currently and find all the misplaced carriables.
Specifically, Due to the token limitation of the LLM, we cannot input all the carriable objects at once for scoring, so we score carriables one by one. For each carriable object, we collect the description of the carriable object, the context-induced affordance of the receptacle it is currently located, and our task instruction, which is ”tide the house.” With the aboved collected information as well as a well-designed prompt, LLM will analyze the current placement of the given carriable object to determine the suitability of the placement, and assign a score to this relevance.
Note that scoring objects individually can lead to variability in standards due to the LLM’s lack of criteria context. To mitigate this, we set a fixed standard in the prompt for the LLM to refer to when rating: 100 points: the placement perfectly meets the task requirements; 0 points: the placement contradicts the task requirements and might negatively impact the task, and thus need to be rearranged; 50 points: It is difficult to judge whether rearranging this object is related to the task. Figure 5 shows an example of the scoring process. We set 50 points as the threshold and all those carriable objects with score no more than 50 are considered to be misplaced and should be rearranged. These carriables are then reorganized in accordance with their respective scores, arranged in ascending order, to facilitate the planning of subsequent rearrangements.
3.3. Object Rearrangement Planning
For each misplaced object, our goal is to find the most suitable receptacle in the scene as the placement target with the context-induced affordance information stored in the AEG. To allow LLM to find the most suitable placement, it is necessary to consider all the relevant receptacles for comparison. However, directly inputting the affordances of all possible receptacles in the AEG to LLM will not only flood it with excessive irrelevant information, impairing its reasoning ability, but also makes it lose focus, increasing the likelihood of hallucinations.
Inspired by Retrieval Augmented Generation (RAG) (Gao et al., 2023a), which retrieves the task-related text chunks from a given database to augment LLM’s task-solving ability and prevent hallucination, we construct a database of receptacle affordance and retrieve the top-k task-related receptacles from this database as suitable candidates and send them to LLM to select the best target, as shown in Figure 6. In this way, the task-irrelevant receptacles as well as the distracting information will be filtered during the retrieving process, and with the rest top-k candidates as prompt, LLM can make accurate placement decision and generate high-performance rearrangement planning.
Receptacle candidate retrieval.
Different from the classic RAG, which splits a long text into multiple chunks to create the database and retrieves the most relevant chunks based on semantic similarity to the query, our dataset consists of descriptions of all receptacle nodes in the AEG, and the top-k most relevant receptacles are retrieved based on task relevance to the query.
To achieve this, we adopt our LLM scorer from misplacement detector with a minor change of prompt and use it to rate the appropriateness of placement based on task instruction. The inputs are same as the misplacement detector, consisting of the task instruction “tide the house”, the carriable object description, as well as the receptacle’s context-induced affordance. However, the scoring standards in the prompt are adjusted accordingly as follow: “ 100 points: The LLM believes this receptacle is the most suitable place in the entire house to place the carriable in order to complete the task; 0 points: The LLM believes that, from the task’s perspective, placing the given carriable in this receptacle should not be considered under any circumstances; 50 points: The LLM believes that placing the object in this receptacle could only be reasonable under certain special conditions. ” With all receptacle rated, we can choose the top k receptacles with the highest scores as candidates and use their context-induced affordance as retrievals for the placement decision generation.
Placement decision generation.
We introduce a LLM-based placement decision generator to select the best receptacle among the proposed candidates and generate rearrangement plan. With the top-k rated receptacles, we feed their context-induced affordance as well as the query (description of carriable to rearrange together with the task) into the placement decision generator. The LLM then selects the best placement target, outputting the placement decision as following format: “ Best receptacle: the name of the most suitable receptacle among the candidates; Analysis: the reason why it thinks this receptacle is the best one compared to others.
With the generated placement target, we can pick up all the misplaced carriable objects and transfer them to their target placement location, completing the rearrangement process.
4. Experiment and Results
Table 1. Comparison to Tidybot on its benchmark by sorting criteria. We calculated the average success rate of different sorting criteria by weighting the success rate with the number of objects within that criteria.
Method Category Attribute Function Subcategory Multiple Average
Tidybot 91.0% 85.6% 93.9% 90.1% 93.5% 90.7%
Ours 91.3% 88.8% 93.7% 89.1% 96.0% 91.4%
Table 2. Comparison with Housekeep (Kant et al., 2022) on Rearrangement and Misplacement detection task. The random selection strategy is added in the comparison as a baseline to demonstrate improvement given by the different methods. The best results for each indicator are highlighted in bold.
Method Rearrangement (Top k NDCG) Misplacement detection
1 2 3 4 5 6 7 8 Accuracy Recall Precision F1 score
Random 0.181 0.193 0.204 0.216 0.234 0.253 0.274 0.295 0.57 0.594 0.815 0.687
HouseKeep (Kant et al., 2022) 0.265 0.249 0.268 0.286 0.302 0.316 0.333 0.351 0.637 0.707 0.831 0.764
Ours 0.648 0.627 0.621 0.635 0.652 0.667 0.684 0.696 0.829 0.88 0.908 0.894
4.1. Benchmark
Datasets.
Two benchmarks were used for testing, one proposed by Tidybot (Wu et al., 2023) and another dataset annotated by us based on the Habitat Synthetic Scenes Dataset (HSSD 200) (Khanna et al., 2023). It is worth noting that the scenes in the Tidybot dataset are all single-room, with generally no more than 5 objects per scene and relatively sparse context information. Therefore, to fully validate the performance of methods in scenarios closer to real-world settings, we established a context-oriented dataset consisting of 8 scenes based on existing complex scene datasets, totaling 36 rooms, 105 areas, and 188 receptacle objects. We randomly place carriable objects into the benchmark scenarios to generate more than 4000 messy scenes for misplacement detection evaluation.
Metrics.
For the household rearrangement task, we evaluated our method mainly in two aspects: misplacement detection and carriable object rearrangement planning. When the ground-truth object placements are available, we measure the detection performance based on accuracy, recall, precision, and F1 score following most existing works (Padilla et al., 2020). To evaluate the performance of carriable object rearrangement planning, we take the best 8 placements for each carriable object to calculate the Normalized Discounted Cumulative Gain (NDCG) based on the labeled ground-truth receptacles. More details about the benchmark and evaluation metrics can be found in the supplementary material.
4.2. Quantitative Comparisons
Comparison on Tidybot benchmark.
We first compared our method with the Tidybot (Wu et al., 2023) on its benchmark for the rearrangement planning task. Note that the Tidybot needs to imitate the given preferences to perform the planning. Two demonstration placements are provided for each receptacle for imitation in the Tidybot’s test. On the other hand, we only take the given demonstration placements as a regular context for affordance analysis during testing. As shown in Table 1, our method can still achieve better performance than Tidybot under this unfair setting, which demonstrates that our AEG can support LLMs in extracting effective information even when the context information is relatively sparse.
Comparison on our context-oriented benchmark.
Note that Tidybot fails when testing on our context-oriented benchmark due to its limited performance in large-scale scenes and the setting where no certain demonstration placements are provided. Thus, we compare our method with a state-of-the-art alternative Housekeep (Kant et al., 2022), a learning-based rearranging method that learns the knowledge priors from the semantic context of the existing scenes. As shown in the comparison results in Table 2, our approach significantly outperforms Housekeep in both placement rearrangement and misplacement detection tasks. It demonstrates that the proposed AEG can significantly better enhance the agent’s understanding of scenes by leveraging additional context such as spatial, geometric, and visual cues besides semantic relationships.
4.3. Ablation Studies
Table 3. Ablation study on different context-induced types via evaluating the misplacement detection performance on our context-oriented dataset.
Context-induced types accuracy recall precision F1
Visual Textual Global
0.774 0.813 0.901 0.852
0.811 0.856 0.907 0.880
0.821 0.878 0.899 0.888
0.828 0.878 0.909 0.892
0.829 0.880 0.908 0.893
Necessity of context-induced affordance.
The key to the success of our method lies in the affordance-enhanced graph (AEG) derived via context-induced affordance analysis. To verify the necessity of AEG with full context information, we compare it with other alternatives using a vanilla scene graph (no local context analysis), a visual-context-enhanced scene graph (only keyframe context is given), or a textual-context-enhanced scene graph (only textual description is given) on our benchmark. The results are presented in Table 3 and Figure 7 (left). We can see that the vanilla scene graph can not provide enough support for both tasks since it can not reflect common-sense knowledge and personalized preference in the scene. Meanwhile, we find that the visual context is more effective for the misplacement detection and the textual context is more essential for the rearrangement. Our method takes both into the affordance analysis and achieves the best performance in both tasks.
Figure 7. Ablation study on context-induced affordance enhancement and LLM-based scorer via measuring the NDCG performance.
Necessity of LLM-based scorer.
Compared to asking the LLM to predict the final receptacle directly based on AEG, we find that adopting an LLM-based scorer to retrieve candidates for an RAG query can provide more reasonable results. To make a fair evaluation, we compare our LLM-based scorer strategy with a similarity-based candidate retrieval method (Gao et al., 2023b) that is widely used in NLP applications. We also added a random selection strategy that serves as a baseline. As presented in the right of Figure 7, our LLM-based scorer can provide significantly better performance than other alternatives. We can effectively filter out a large number of irrelevant options with the LLM-based scoring, which greatly reduces the difficulty for the LLM in making the final decision.
4.4. Qualitative Results
To better illustrate the pipeline of our method, we show some examples of context-induced affordance enhancement in Figure 8. Note how the agent derives more accurate functional descriptions of target objects than the original general semantic tags. Besides, we showcase several examples of carriable object placement planning based on AEG in Figure 9 with real-world scenarios and Figure 10 with synthetic environments. We can see that the agent can plan object placements that better fit the scene context based on AEG. More results can be found in supplementary material.
5. Conclusion
We introduced a method to automatically mine object functionalities aligned with user preferences from a scene by using multi-modal foundation models prompted with scene graphs and RGB keyframes. The resulting affordance-enhanced graph effectively detects misplaced objects and determines their proper locations.
Limitation and future work.
Our affordance enhancement, which integrates contextual information, is contingent upon the accuracy of the initial scene graph, thus it cannot rectify errors inherent in its construction. Moving forward, we are keen on pursuing the joint optimization of the initial scene graph through end-to-end and/or active learning approaches. The pivotal role of RGB images in affordance analysis necessitates careful keyframe selection to prevent misleading contexts caused by the occlusion of pertinent objects or the visibility of extraneous ones; one failure case with inappropriate keyframe selection in the latter case is shown in the supplementary material. Consequently, it is a compelling area of study to explore optimal keyframe selection techniques that reduce ambiguities in image prompting.